A sample of oxygen gas collected at a pressure of 0.905 atm and a temperature of 278 K is found to occupy a volume of 651 milliliters. How many moles of O2 gas are in the sample?
Answer #1

Criven - P- 0.905 atm = 2785 - 65ml = 0.6512 n=2 dince, pressure is in ahm & Volume in litere, R = 0.0821. Lamm k mol! PV=nRT